OpenStudy (anonymous):

Change the following equations to y=mx+b form so I can graph. 3x+2y=-8 2x-3y=-1

OpenStudy (anonymous):

Change the equations to y = mx + b, then graph.

OpenStudy (anonymous):

Where they intersect is the answer.

OpenStudy (anonymous):

Thats what I need help with. Changing it.

OpenStudy (anonymous):

What? Thats not y=mx+b form.

OpenStudy (anonymous):

Start by solving one equation for y and the other equation for x

OpenStudy (anonymous):

Sammi thats solving it not changing the form.

OpenStudy (anonymous):

So Joe the first one would be x=-2/3y-2 2/3

OpenStudy (anonymous):

right. now you can substitute this into the second equation. so, instead of x, use (-2/3y-2 2/3)